OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 15 meters and legs that measure 9 meters and 12 meters. What is the surface area of this prism if the height of the prism is 11 meters? 450 m2 486 m2 504 m2 612 m2

OpenStudy (anonymous):

Calculate the lateral areas (the sides are rectangles, so it's no big deal) then add twice the area of the base. Since it is a right triangle, the height and base in the area formula will be the two legs.

OpenStudy (anonymous):

If you do a little algebra along the way, it makes it even easier.\[A=ah+bh+ch+2(\frac{ab}{2})=h(a+b+c)+ab=11(9+12+15)+9(12)=504\]

OpenStudy (anonymous):

so its the third one?

OpenStudy (anonymous):

last time they told me it was the last one and it was wrong

OpenStudy (anonymous):

Look at my math. What do you think?

OpenStudy (anonymous):

ur math dosent apear

OpenStudy (anonymous):

now it does

OpenStudy (anonymous):

its 396? butt it dosent match

OpenStudy (anonymous):

answer????

OpenStudy (anonymous):

Assuming that the prism is a right triangular base prism, its surface area is the sum of its lateral area and the area of its two triangular bases. SA = ph + 2 B where p is the perimeter of the base and B is the area of each base. SA = ( 9 + 12 + 15) * 11 + 2 (1/2) (9)(12) SA = 36 * 11 + 2 (108) = 612 square meters

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

lawrenceh123: pretty sure your next-to-last phrase and your total are wrong. The two and the one half cancel, so it ends up 396 plus 108 is 504.

OpenStudy (anonymous):

true i got the answer wrong thanks

OpenStudy (anonymous):

So were you right or me, im confused